An annuity immediate pays $500 per month for the first three years. After that the annuity payments increase by $50 per month for five years and then remain level for an additional six years. At a nominal rate of annual interest of 12% convertible monthly what is the present value of this annuity?

The answer in the back is: 140339.571
Answer #1
.A. $500 per month for first three years:
Present Value (PV) of payment=(Payment)/((1+i)^N)
i=monthly interest=(12/12)%=1%=0.01
N=Month of payment
Payment and PV of payments are given below:
N A B=A/(1.01^N)
Month Payment PV of payment
0 $500 500
1 $500 495.049505
2 $500 490.148025
3 $500 485.295074
4 $500 480.490172
5 $500 475.732844
6 $500 471.022618
7 $500 466.359027
8 $500 461.741611
9 $500 457.169912
10 $500 452.643477
11 $500 448.161859
12 $500 443.724613
13 $500 439.3313
14 $500 434.981485
15 $500 430.674737
16 $500 426.410631
17 $500 422.188744
18 $500 418.008657
19 $500 413.869958
20 $500 409.772235
21 $500 405.715084
22 $500 401.698103
23 $500 397.720894
24 $500 393.783064
25 $500 389.884221
26 $500 386.023982
27 $500 382.201962
28 $500 378.417784
29 $500 374.671073
30 $500 370.961459
31 $500 367.288573
32 $500 363.652053
33 $500 360.051537
34 $500 356.486671
35 $500 352.9571
TOTAL $ 15,204.29
.B. Increasing by $50 per month for next five years
N A B=A/(1.01^N)
Month Payment PV of payment
36 $550 384.408722
37 $600 415.20294
38 $650 445.349689
39 $700 474.858769
40 $750 503.739854
41 $800 532.002486
42 $850 559.656081
43 $900 586.709927
44 $950 613.173192
45 $1,000 639.054916
46 $1,050 664.364021
47 $1,100 689.10931
48 $1,150 713.299466
49 $1,200 736.943056
50 $1,250 760.048531
51 $1,300 782.62423
52 $1,350 804.678378
53 $1,400 826.21909
54 $1,450 847.254371
55 $1,500 867.792118
56 $1,550 887.84012
57 $1,600 907.406064
58 $1,650 926.497528
59 $1,700 945.121991
60 $1,750 963.286828
61 $1,800 980.999316
62 $1,850 998.26663
63 $1,900 1015.09585
64 $1,950 1031.49396
65 $2,000 1047.46784
66 $2,050 1063.0243
67 $2,100 1078.17002
68 $2,150 1092.91162
69 $2,200 1107.25561
70 $2,250 1121.20843
71 $2,300 1134.77641
72 $2,350 1147.9658
73 $2,400 1160.78278
74 $2,450 1173.23342
75 $2,500 1185.32372
76 $2,550 1197.0596
77 $2,600 1208.44689
78 $2,650 1219.49134
79 $2,700 1230.19862
80 $2,750 1240.57433
81 $2,800 1250.62399
82 $2,850 1260.35303
83 $2,900 1269.76682
84 $2,950 1278.87064
85 $3,000 1287.66972
86 $3,050 1296.16919
87 $3,100 1304.37412
88 $3,150 1312.28952
89 $3,200 1319.92031
90 $3,250 1327.27135
91 $3,300 1334.34744
92 $3,350 1341.15329
93 $3,400 1347.69357
94 $3,450 1353.97286
95 $3,500 1359.9957
TOTAL $ 59,554.86
.C. $3500 per month for next six years
N A B=A/(1.01^N)
Month Payment PV of payment
96 $3,500 1346.5304
97 $3,500 1333.19841
98 $3,500 1319.99843
99 $3,500 1306.92914
100 $3,500 1293.98924
101 $3,500 1281.17747
102 $3,500 1268.49254
103 $3,500 1255.93321
104 $3,500 1243.49823
105 $3,500 1231.18636
106 $3,500 1218.9964
107 $3,500 1206.92713
108 $3,500 1194.97736
109 $3,500 1183.1459
110 $3,500 1171.43158
111 $3,500 1159.83325
112 $3,500 1148.34975
113 $3,500 1136.97995
114 $3,500 1125.72272
115 $3,500 1114.57696
116 $3,500 1103.54154
117 $3,500 1092.61539
118 $3,500 1081.79741
119 $3,500 1071.08655
120 $3,500 1060.48173
121 $3,500 1049.98191
122 $3,500 1039.58605
123 $3,500 1029.29312
124 $3,500 1019.1021
125 $3,500 1009.01198
126 $3,500 999.02176
127 $3,500 989.130455
128 $3,500 979.337084
129 $3,500 969.640678
130 $3,500 960.040275
131 $3,500 950.534926
132 $3,500 941.123689
133 $3,500 931.805632
134 $3,500 922.579834
135 $3,500 913.44538
136 $3,500 904.401367
137 $3,500 895.446898
138 $3,500 886.581087
139 $3,500 877.803056
140 $3,500 869.111937
141 $3,500 860.506868
142 $3,500 851.986998
143 $3,500 843.551483
144 $3,500 835.199488
145 $3,500 826.930187
146 $3,500 818.742759
147 $3,500 810.636395
148 $3,500 802.610292
149 $3,500 794.663656
150 $3,500 786.795699
151 $3,500 779.005642
152 $3,500 771.292715
153 $3,500 763.656153
154 $3,500 756.095201
155 $3,500 748.60911
156 $3,500 741.197139
157 $3,500 733.858553
158 $3,500 726.592627
159 $3,500 719.398641
160 $3,500 712.275882
161 $3,500 705.223645
162 $3,500 698.241233
163 $3,500 691.327954
164 $3,500 684.483122
165 $3,500 677.706062
166 $3,500 670.996101
167 $3,500 664.352575
TOTAL $ 69,564.31
A PV of $500 per month for first three years: $    15,204.29
B PV of payment Increasing by $50 per month for next five years $    59,554.86
C PV of $3500 per month for next six years $    69,564.31
Total Present Value of Payments for(3+5+6)=14 years $ 144,323.46